Chuck Wagon Grills, Inc., makes a single product—a handmade specialty barbecue grill that it sells for $200. Data for last year’s operations follow:

  

Units in beginning inventory 0
Units produced 10,400
Units sold 8,800
Units in ending inventory 1,600
Variable costs per unit:
Direct materials $ 60
Direct labor 20
Variable manufacturing overhead 10
Variable selling and administrative 30
Total variable cost per unit $ 120
Fixed costs:
Fixed manufacturing overhead $ 150,000
Fixed selling and administrative 290,000
Total fixed costs $ 440,000

Required:

1. Assume that the company uses variable costing. Compute the unit product cost for one barbecue grill.

2. Assume that the company uses variable costing. Prepare a contribution format income statement for last year.

3. What is the company’s break-even point in terms of the number of barbecue grills sold?
Answer #1
1
Direct materials 60
Direct labor 20
Variable manufacturing overhead 10
Unit product cost 90
2
Chuck Wagon Grills, Inc.
Variable Costing Income Statement
Sales 1760000 =8800*200
Variable expenses:
Variable cost of goods sold 792000 =8800*90
Variable selling and administrative 264000 =8800*30
1056000
Contribution margin 704000
Fixed expenses:
Fixed manufacturing overhead 150000
Fixed selling and administrative 290000
440000
Net operating income 264000
3
Total fixed costs 440000
Divide by Unit Contribution margin 80 =200-120
Break-even point in units 5500